OpenStudy (anonymous):

You pick 5 cards out of a deck of 52. What is the probability that you get exactly two spades?

OpenStudy (anonymous):

I tried it and thought it was \[\left(\begin{matrix}5 \\ 2\end{matrix}\right)(\frac{ 1 }{ 4 })^2(\frac{ 3 }{ 4 })^3\] But someone else posted somewhere on the internet that it would be \[\frac{ 13 }{ 52 } * \frac{ 12 }{ 51 } * \frac{ 39 }{ 50 } * \frac{ 38 }{ 49 } * \frac{ 37 }{ 48 }\] But I think now it's: \[\frac{ \left(\begin{matrix}5 \\ 2\end{matrix}\right)(\frac{ 1 }{ 4 })^2(\frac{ 3 }{ 4 })^3 }{ \left(\begin{matrix}52 \\ 5\end{matrix}\right) }\]

OpenStudy (anonymous):

there are 13 spades, and 40 non-spades. so (13C2) (40C3) / (52C5)

OpenStudy (anonymous):

oops, i meant 39 non-spade XD

OpenStudy (anonymous):

(13C2) (39C3) / (52C5)

OpenStudy (anonymous):

Oh that makes sense! thank you! I have another similar question too, if you don't mind

OpenStudy (anonymous):

You pick 5 cards out of a deck of 52. What is the probability that we get at least one card of each suit?

OpenStudy (anonymous):

was the solution given? i can work it out but not sure if it's right

OpenStudy (anonymous):

No:( I wish lol

OpenStudy (anonymous):

we I have (13C1) (13C1) (13C1) (13C1) (48) / (52C5) each of the first 4 cards are guaranteed to in a different suit. The 5th card can be any card that is remaining

OpenStudy (anonymous):

I mean that sounds right to me lol. Thank you!

OpenStudy (anonymous):

I'm not sure it's that, but i'm also not sure it's what I have: \[\frac{ \left(\begin{matrix}4 \\ 1\end{matrix}\right)\left(\begin{matrix}4 \\ 1\end{matrix}\right) \left(\begin{matrix}4 \\ 1\end{matrix}\right) \left(\begin{matrix}4 \\ 1\end{matrix}\right) \left(\begin{matrix}48 \\ 1\end{matrix}\right) }{ \left(\begin{matrix}52 \\ 5\end{matrix}\right) }\] It could be this because you have to choose 1 of 4 suits for each of the first 4 cards, and the last card could be one of 48 other cards that could be picked in any order

OpenStudy (anonymous):

Is that how that works?

OpenStudy (anonymous):

each suit has 13 right?

OpenStudy (anonymous):

13 cards

OpenStudy (anonymous):

Yeah, but you're just picking 1 of 4 suits, so I would think that you would just choose 1 of the 4 suits per card and then the last one is one of the remaining 48 cards but then I got confused because my prof had told me that the n's in the numerator had to add up to equal the n in the denominator (for the n C r) and the same thing for the r's

OpenStudy (anonymous):

And any of the cards can be in any order

OpenStudy (anonymous):

not sure :/